OpenStudy (anonymous):

Solve for w in P = 2 w + 2 l ,P = 38 and l= 12

OpenStudy (anonymous):

So basically: 38=2*W +2*12

OpenStudy (anonymous):

ummm is it not that you need to find w? then move the equation around to find w

OpenStudy (anonymous):

kay.

OpenStudy (anonymous):

How do I do that?

OpenStudy (anonymous):

I'm not very good at doing that...

OpenStudy (anonymous):

well you have 38=2w+24 so move the 24 to the left by eliminating it on the right by -24 meaning also reducing -24 from the left => so 38-24=2w+24-24 that is 14=2w now divide both by two to obtain a 1 before w since 2/2=1 and 1w=w so 14/2=2/2w=w => w=7

OpenStudy (anonymous):

Thanks.